OpenStudy (anonymous):

Find an identity for cos(4t) in terms of cos(t)?

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't get it?

OpenStudy (anonymous):

\[\frac{ -\sin(8t) }{\cos(8t)-1 }\]

OpenStudy (anonymous):

The double angle formula for cos says that\[\cos(2A)=2\cos^2A-1\] so \[\cos(4t)=\cos(2(2t))=2\cos^2(2t)-1\]